CNNBC recently reported that the mean annual cost of auto insurance is 1009 dollars. Assume the standard deviation is 223 dollars. You take a simple random sample of 100 auto insurance policies.

Find the probability that a single randomly selected value is less than 983 dollars.
P(X < 983) =

Find the probability that a sample of size n=100n=100 is randomly selected with a mean less than 983 dollars.
P(¯xx¯ < 983) =

Enter your answers as numbers accurate to 4 decimal places.
Answer #1

solution Given than u = 1009 dollars 0= 223 dollars - P(x<983) 983-1009 1 6 223 Р / = P(24-0.12 ) = 0.4522 I P(x< 983) = 0.45